# MLxOR: Mathematical Foundations and Operational Integration of Machine Learning for Uncertainty-Aware Decision-Making

### v1

@@ Abstract
Aging infrastructure portfolios pose a critical resource allocation challenge: deciding which structures require intervention and which can safely remain in service. Structural assessments must balance the trade-off between cheaper, conservative analysis methods and accurate but costly simulations that do not scale portfolio-wide. We propose Bayesian neural network (BNN) surrogates for rapid structural pre-assessment of worldwide common bridge types, such as reinforced concrete frame bridges. Trained on a large-scale database of non-linear finite element analyses generated via a parametric pipeline and developed based on the Swiss Federal Railway’s bridge portfolio, the models accurately and efficiently estimate high-fidelity structural analysis results by predicting code compliance factors with calibrated epistemic uncertainty. Our BNN surrogate enables fast, uncertainty-aware triage: flagging likely critical structures and providing guidance where refined analysis is pertinent. We demonstrate the framework’s effectiveness in a real-world case study of a railway underpass, showing its potential to significantly reduce costs and emissions by avoiding unnecessary analyses and physical interventions across entire infrastructure portfolios.
@@ 1 Introduction
Current assessment practices follow a tiered, so called levels-of-approximation approach [5]: Starting with simplified, conservative analyses that contain large safety margins to compensate for modeling abstractions. If these methods do not demonstrate compliance suggesting that an intervention is necessary, asset managers can commission more refined simulations such as non-linear finite element analysis (NLFEA). While NLFEA is expensive, time-consuming, and requires detailed inputs and expertise, making it unsuitable for portfolio-wide use, it can uncover structural reserve capacity, avoiding costly strengthening measures and enabling structures to remain in service beyond their original design life. However, the benefit of such detailed analyses is unknown a priori. In some cases, NLFEA merely confirms the findings of conservative methods at a significant extra cost.
@@
Related work Data-driven surrogates have been widely applied to approximate computationally expensive simulations in engineering, including structural health monitoring and capacity estimation [2, 9, 11, 16]. However, particularly for structural engineering most approaches remain deterministic and lack uncertainty quantification. Probabilistic methods such as Gaussian Processes, deep ensembles, and BNNs provide predictive distributions instead of point estimates [6], with recent studies demonstrating their effectiveness for pavement response prediction [1], bridge integrity assessment [13], and materials modeling [14].
@@
Orange
$\mu>1$ and
$\mu-2\kappa\sigma<1$
Likely safe, but uncertainty is high. A refined analysis is recommended, especially if the predicted compliance factor is only slightly above 1.
Green
$\mu>1$ and
$\mu-2\kappa\sigma\geq 1$
High confidence in compliance. Not to prioritize. No further analysis is required at this stage. Simplified analysis methods are predicted to be sufficient.
@@ 2 Methods
Data generation with parametric NLFEA. To obtain training for supervised learning, we built an automated pipeline that generates models of reinforced concrete frame bridges from a parameter vector, evaluates them with NLFEA, and extracts code-compliance labels. The pipeline links the programs Rhino/Grasshopper (for geometry generation) with Ansys Mechanical APDL (for NLFEA), where the Cracked Membrane Model [7, 15] is implemented as a layered shell element user material. This orchestration is automated via the StrucEngLib plug-in [8]. For each simulation, the pipeline validates structural safety according to Swiss design codes and outputs three global compliance factors $\boldsymbol{\eta}(\mathbf{x})=$$(\eta_{M,c,\min},\eta_{M,s,\min},\eta_{V,\min})$. These quantify the structural safety margin in bending (for concrete and reinforcing steel) and shear, computed as the minimum resistance-to-demand ratios across all elements, with $\eta\geq 1$ indicating code compliance. Since each NLFEA run is expensive ($\approx\!25$ minutes per bridge, $\approx\!2.5$ minutes with 10-way parallelism) [^1] Note that in practice, the high cost of NLFEA is caused less by the simulation time itself and more by the manual effort required to build a validated digital model of a specific bridge., we adopt a sampling strategy that balances coverage and efficiency. To ensure realistic variability, parameter ranges were derived from a database analysis of the Swiss Federal Railways (SBB) bridge inventory, with expert feedback. Latin Hypercube Sampling (LHS) is used to broadly cover the input space, followed by adaptive resampling based on kernel density estimation to focus simulations near the safety-critical region around $\eta\approx 1$ (cf. Table 3 in [10]). The resulting dataset of $\approx\!11$k spans common structures in the SBB portfolio while providing high resolution near the decision boundary.
@@ 3 Results
Decision-focused accuracy. Prediction accuracy and uncertainty estimation are strongest in the safety-critical range $0.5\leq\eta\leq 1.5$ (Figure 2). Predictive intervals narrow near $\eta\approx 1$ and widen in uncritical regions, due to our targeted resampling and dedicated loss. Among the targets, $\eta_{M,s,\min}$ is predicted most reliably (RMSE 0.10, MAPE 4.8%), $\eta_{M,c,\min}$ shows moderate error, and $\eta_{V,\min}$ remains most challenging due to the fact that shear is governed by abrupt local effects (cf. Table 2).
@@
Table 2: Predictive performance in the safety-critical region $\eta\in[0.5,1.5]$. CB reported after post hoc scaling.
@@
Calibration of predictive uncertainty. After training, the predictive intervals computed as $\mu\pm z(p)\sigma$, where $z(p)$ denotes the standard normal quantile for a given confidence level $p$, indicate that the BNN outputs are mildly overconfident in their uncertainty estimates. These intervals capture fewer true values than expected, causing calibration curves to fall below the diagonal and CB to be positive. We therefore apply a post-hoc scale factor per head (chosen on the validation set: $\kappa=1.3$ for steel, $1.7$ for concrete, $2.8$ for shear) to align nominal and empirical coverage, reduce TCE, and obtain mildly under-confident intervals (CB <0<0, Table 2) on the held out data, which is desirable in this safety-critical setting. The Gaussian predictive distribution assumption holds well overall, but less so for shear, due to its brittle and locally governed behavior.
@@ 4 Conclusion
We present a scalable, uncertainty-aware surrogate for structural pre-assessment of common bridge types. Trained on high-fidelity NLFEA data, our BNN model predicts code compliance with sufficient accuracy and calibrated uncertainty, supporting crucial triage decisions and whether refined analysis is warranted. Demonstrated on a real-world case study, the approach enables portfolio-scale screening even from limited input detail. It helps reduce unnecessary assessments and focuses scarce resources where they are most impactful, a critical need given today’s aging infrastructure portfolios.
@@
We acknowledge support and data access from Swiss Federal Railways (SBB). Parts of this work were conducted at ETH Zurich and TU Darmstadt. See journal version for extended acknowledgments.

### v2

@@ Abstract
Aging infrastructure portfolios pose a critical resource allocation challenge: deciding which structures require intervention and which can safely remain in service. Today’s structural assessment approaches do not scale to portfolio level, as they require time-consuming manual digital modeling and computationally expensive simulations for each structure. We propose Bayesian neural network (BNN) surrogates for rapid structural pre-assessment of worldwide common bridge types, such as reinforced concrete frame bridges. Trained on a large-scale database of non-linear finite element analyses generated via a parametric pipeline and developed based on the Swiss Federal Railway’s bridge portfolio, the models accurately and efficiently estimate high-fidelity structural analysis results by predicting code compliance factors with calibrated epistemic uncertainty. Our BNN surrogate enables fast, uncertainty-aware triage: flagging likely critical structures and providing guidance where refined analysis is pertinent. We demonstrate the framework’s effectiveness in a real-world case study of a railway underpass, showing its potential to significantly reduce costs and emissions by avoiding unnecessary analyses and physical interventions across entire infrastructure portfolios.
@@ 1 Introduction
Current assessment practices follow a tiered, so called levels-of-approximation approach [5]: Starting with simplified, conservative analyses that contain large safety margins to compensate for modeling abstractions. If these methods do not demonstrate compliance suggesting that an intervention is necessary, asset managers can commission more refined simulations such as non-linear finite element analysis (NLFEA). While NLFEA is expensive, time-consuming, and requires detailed inputs and expertise, it can uncover structural reserve capacity, avoiding costly strengthening measures and enabling structures to remain in service beyond their original design life. However, the benefit of such detailed analyses is unknown a priori. In some cases, NLFEA merely confirms the findings of conservative methods at a significant extra cost. Furthermore, the typically high manual modeling and computational effort involved makes current assessment workflows unscalable to portfolio level.
@@
Data-driven surrogates have been widely applied to approximate computationally expensive simulations in engineering, including structural health monitoring and capacity estimation [2, 9, 11, 17]. However, for structural engineering most approaches remain deterministic and lack uncertainty quantification. Probabilistic methods such as Gaussian Processes, deep ensembles, and BNNs provide predictive distributions instead of point estimates [6], with recent studies demonstrating their effectiveness for pavement response prediction [1], bridge integrity assessment [14], and materials modeling [15].
@@
Orange $\mu>1$ and $\mu-2\kappa\sigma<1$ Likely safe, but uncertainty is high. A refined analysis is recommended, especially if the predicted compliance factor is only slightly above 1.
Green $\mu>1$ and $\mu-2\kappa\sigma\geq 1$ High confidence in compliance. Not to prioritize. Simplified analysis methods are predicted to be sufficient.
@@ 2 Methods
Data generation with parametric NLFEA. To obtain training data for supervised learning, we built an automated pipeline that generates models of reinforced concrete frame bridges from a parameter vector, evaluates them with NLFEA, and extracts code-compliance labels. The pipeline links the programs Rhino/Grasshopper (for geometry generation) with Ansys Mechanical APDL (for NLFEA), where the Cracked Membrane Model [7, 16] is implemented as a layered shell element user material. This orchestration is automated via the StrucEngLib plug-in [8]. For each simulation, the pipeline validates structural safety according to Swiss design codes and outputs three global compliance factors $\boldsymbol{\eta}(\mathbf{x})=$$(\eta_{M,c,\min},\eta_{M,s,\min},\eta_{V,\min})$. These quantify the structural safety margin in bending (for concrete and reinforcing steel) and shear, computed as the minimum resistance-to-demand ratios across all elements, with $\eta\geq 1$ indicating code compliance. Since each NLFEA run is expensive ($\approx\!25$ minutes per bridge, $\approx\!2.5$ minutes with 10-way parallelism) [^1] Note that in practice, the high cost of NLFEA is caused less by the simulation time itself and more by the manual effort required to build a validated digital model of a specific bridge., we adopt a sampling strategy that balances coverage and efficiency. To ensure realistic variability, parameter ranges were derived from a database analysis of the Swiss Federal Railways (SBB) bridge inventory, with expert feedback. Latin Hypercube Sampling (LHS) is used to broadly cover the input space, followed by adaptive resampling based on kernel density estimation to focus simulations near the safety-critical region around $\eta\approx 1$ (cf. Table 3 in [10]). The resulting dataset of $\approx\!11$k samples spans common structures in the SBB portfolio while providing high resolution near the decision boundary.
@@ 3 Results
Decision-focused accuracy. Prediction accuracy and uncertainty estimation are strongest in the safety-critical range $0.5\leq\eta\leq 1.5$ (Fig. 2). Predictive intervals narrow near $\eta\approx 1$ and widen in uncritical regions, due to our targeted resampling and dedicated loss. Among the targets, $\eta_{M,s,\min}$ is predicted most reliably (RMSE 0.10, MAPE 4.8%), $\eta_{M,c,\min}$ shows moderate error, and $\eta_{V,\min}$ remains most challenging due to the fact that shear is governed by abrupt local effects (cf. Table 2).
@@
Table 2: Predictive performance in the safety-critical region $\eta\in[0.5,1.5]$. CB reported after post-hoc scaling.
@@
Calibration of predictive uncertainty. After training, the predictive intervals computed as $\mu\pm z(p)\sigma$, where $z(p)$ denotes the standard normal quantile for a given confidence level $p$, indicate that the BNN outputs are mildly overconfident in their uncertainty estimates. These intervals capture fewer true values than expected, causing calibration curves to fall below the diagonal and CB to be positive. We therefore apply a post-hoc scale factor per head (chosen on the validation set: $\kappa=1.3$ for steel, $1.7$ for concrete, $2.8$ for shear) to align nominal and empirical coverage, reduce TCE, and obtain mildly under-confident intervals (CB <0<0, Table 2) on the held-out data, which is desirable in this safety-critical setting. The Gaussian predictive distribution assumption holds well overall, but less so for shear, due to its brittle and locally governed behavior.
@@ 4 Conclusion
We present a scalable, uncertainty-aware surrogate for structural pre-assessment of common bridge types. Trained on high-fidelity NLFEA data, our BNN model efficiently predicts code compliance based on a few structural input parameters with sufficient accuracy and calibrated uncertainty, supporting crucial triage decisions and whether refined analysis is warranted. Demonstrated on a real-world case study, the approach enables portfolio-scale screening even from limited input detail. It helps reduce unnecessary assessments and focuses scarce resources where they are most impactful, a critical need given today’s aging infrastructure portfolios.
@@
We acknowledge support and data access from Swiss Federal Railways (SBB). See journal publication [10] for extended acknowledgments.